# Verification and Validation for Trustworthy Scientific Machine Learning

### v1

@@ 1 Introduction
Scientific machine learning (SciML) integrates machine learning (ML) into scientific workflows to enhance system simulation and analysis, with an emphasis on computational modeling of physical systems. This field emerged from Department of Energy workshops and initiatives starting in 2018, which also identified the need to increase “the scale, rigor, robustness, and reliability of SciML necessary for routine use in science and engineering applications” [5]. The field’s subsequent growth through funding initiatives, conference themes, and high-profile publications stems from its ability to unite ML’s predictive power with the domain knowledge and mathematical rigor of computational science and engineering (CSE). However, this surge in SciML development has outpaced good practices and reporting standards for building trust [66, 51, 109, 117].
@@
Rigorous modeling practices require sustained effort from the scientific community. CSE has established standards over decades for model reporting and assessment, including verification and validation, across biology [80], earth sciences [79, 45], and engineering [95, 84, 89, 1]. These standards help users and decision-makers document limitations, uncertainties, and modeling choices to prevent resource misallocation and maintain model credibility. Similar SciML efforts have only recently begun [51, 109, 117]. Developers and users of SciML must establish best practices for the development, reporting, and evaluation of new SciML paradigms.
@@ 2.1 Scope
This paper focuses on SciML that enhances computational modeling and simulation of physical systems. Specifically, we address SciML approaches that support computational workflows where a model must be evaluated repeatedly within larger algorithmic procedures. In scientific computing, these are known as “outer-loop” processes because they form an additional computational layer that wraps around the core model solution. Such processes include: (1) explanatory modeling, which searches for insights into complex system behaviors; (2) uncertainty quantification, which infers model parameters from observations and propagates them to predictive distributions [28]; (3) sensitivity analysis, which identifies factors with greatest influence on predictions [88]; (4) experimental design, which predicts which data collections maximize understanding or decision value [87]; and (5) optimal design and control, which optimizes objective functions by adjusting decision variables under constraints [65, Ch. 1]. To support these outer-loop processes, predictive SciML encompasses four main classes of methods, each presenting different computational challenges: surrogate modeling, model discovery, hybrid CSE-SciML modeling, and outer-loop learning.
Surrogate modeling builds low-cost approximations of computationally expensive models for outer-loop analyses. Surrogate models can map model inputs to specific quantities of interest [22, 103, 10]. Surrogate models can also approximate complete solutions to governing equations via data-driven reduced-order modeling [81, 91, 12], operator learning with Gaussian processes [69] and neural networks [39, 54, 63, 8, 58, 76, 27], and physics-informed neural networks (PINNs) [114, 82, 30], which encode physical constraints directly into the neural network training. Finally, surrogate models can provide solutions to differential equations without simulation data [99, 24].
@@
Outer-loop learning leverages ML to predict solutions under varying conditions. Applications include surrogate models for expensive forward simulations in PDE-constrained optimization [110, 64] and Bayesian inverse problems [14]. Reinforcement learning solves PDE-constrained optimization across different conditions [101]. ML methods can also learn solutions to these problems directly. For example, ML models learn Bayesian posterior distributions as marginals of joint distributions for fast online inference with arbitrary observational data [6, 7, 111].
@@ 2.2 Comparing CSE and SciML model development and deployment
Making scientific claims using SciML models involves two steps: model development and model deployment. While SciML models can complement CSE models, their development processes differ, though deployment processes remain similar, as shown in Figure 1. To illustrate these steps, we use an example quantifying uncertainty in a hybrid SciML-CSE ice-sheet model of the Humboldt Glacier in Greenland, with unknown glacier-landmass friction. Our example is loosely based on the study by He et al. [37], which exemplifies thorough reporting of a SciML model’s strengths and weaknesses.
@@
CSE model development begins with physics-based governing equations. In our example, while nonlinear Stokes equations accurately model shear-thinning behavior for ice sheets, simpler models like shallow-shelf approximation (SSA) are often used for computational tractability [37]. The second step implements numerical methods to solve the governing equations while respecting properties like mass and energy conservation. This involves choosing a fixed ansatz (e.g., a finite element basis) and applying numerical procedures using linear algebra, Newton iteration, and time-stepping. For instance, He and coworkers [37] used continuous piecewise linear elements for thickness and velocity fields with Newton iteration and semi-implicit time-stepping.
@@
In contrast to CSE, typical SciML development relies on training procedures that require observational data or the ability to query an existing numerical model. Consequently, as shown in Figure 1, the development begins with collecting input-output training data. For example, He et al. [37] built a DeepONet [63] to map ice-sheet thickness and friction to glacier velocity, using data from an SSA-based CSE model to accelerate glacier evolution and uncertainty quantification. The second step sets model hyperparameters (e.g., neural-network width and depth). The third optimizes model parameters (e.g., neural-network weights and biases) by minimizing the difference between model outputs and training data. He et al. [37] used Adam optimization to minimize the mean-squared error between predicted and training velocities. Often, hyperparameters must be tuned through formal optimization or informal assessment of model accuracy across different settings. Note that some SciML models [99, 24] combine CSE and SciML development steps shown in Figure 1, training without data by minimizing objectives that enforce known governing equations.
@@
The development and deployment of SciML models for specific scientific applications follows distinct phases. When using these models to make scientific claims—–like predicting the Humboldt glacier’s impact on sea level in the future—–deployment involves three key steps beyond the initial model development. First, we collect calibration data representing the system’s current state (distinct from training data, which spans multiple states). Second, we calibrate the model through optimization to match this system-specific data. In our ice-sheet example, this means solving an inverse problem for friction parameters to match recent velocity measurements [46], notably optimizing model inputs rather than SciML parameters like neural-network weights. Finally, we run the model under various future conditions or plausible friction fields, such as those obtained from Bayesian calibration.
@@ 2.3 Leveraging an existing CSE framework for trustworthy SciML
As highlighted by Broderick et al. [11], trust in SciML and CSE models can be gained or lost at any step of development and deployment (Figure 1). Building trust requires demonstrating that the mathematical model adequately meets analysis goals despite being an approximation. The model must then be verified by comparing theoretical and empirical results to ensure it solves its mathematical formulation correctly. Additionally, rigorous code testing must be implemented to reduce software bugs, and data provenance must be documented to ensure relevance for future conclusions and decisions.
@@
While V&V practices are well-established in scientific computing, the emergence of SciML brings new concerns that existing guidelines don’t fully address. SciML combines aspects of both scientific computing and ML, presenting unique verification and validation challenges beyond either field. We argue that adapting and extending proven V&V principles from scientific computing can provide a foundation for building trustworthy SciML models. To this end, we make recommendations in four areas: problem definition, verification, validation, and continuous credibility building (see Figure 2). Problem definition and verification (Sections 3 and 4) expand on the model development steps in Figure 1, while validation and reporting (Sections 5 and 6) detail the deployment steps. Validation is generally most pertinent only to papers that utilize SciML to support a scientific claim for a specific application. However, all other components illustrated in Figure 1 are applicable to all SciML papers, including those that present algorithmic advances, which are relevant to a broad spectrum of scientific applications.
Like Broderick et al. [11], we view trustworthiness as a spectrum rather than binary. Following Oberkampf’s predictive capability maturity model [75], the four components in Figure 2 build confidence through continuous testing and refinement as new data emerge. Each study should mature progressively in its development process through these four areas. The following sections identify opportunities for improving SciML practice, along with corresponding recommendations.
We do not claim that these four components or our recommendations are comprehensive or universally applicable. Instead, we aim to stimulate community efforts to develop SciML guidelines and standards. We hope to advance practice beyond the basic implementation of our recommendations toward more thorough, deliberate, and well-documented approaches to SciML development and deployment.
@@ 3.1 Model purpose
A SciML model’s purpose, as discussed in Section 2.1, dictates all subsequent modeling choices. This purpose determines required outer-loop processes and essential properties.
To highlight the importance of specifying the target outer-loop process, consider a model used for explanatory modeling. An explanatory model must simulate all system processes, like ice-sheet thickness and velocity evolution. In contrast, a risk assessment model needs only decision-relevant quantities, like ice-sheet mass loss under varying emissions scenarios. Design and control models, meanwhile, have different requirements than those for risk assessment. The model purpose dictates the data types and formulations needed to train a SciML model. The impact of this purpose on data requirements and physics-informed objectives varies by application domain. Thus, the exact model formulation should be chosen in light of these problem-specific considerations.
@@ 3.2 Verification, calibration, validation and application domain
The trustworthy development and deployment of a model (see Figure 2) requires using the model in regards to verification, calibration, validation, and application domains. These domains are defined by the conditions under which the model operates during these respective phases and must be defined before model construction because they determine viable model classes. Key features include boundary conditions, forcing functions, geometry, and timescales. For ice sheets, examples include surface mass balance, land mass topography, and ocean temperatures.
Each domain will often require the prediction of different quantities of interest under different conditions. Moreover the complexity of the processes being modeled typically increase when transitioning from verification to calibration, to validation to prediction. Additionally the amount of data to complement or inform the model decreases as we move through these domains. For example, the verification domain for our ice-sheet examplar predicts the entire state of the ice-sheet for simple manufactured or analytical solutions. The calibration domain predicts Humboldt Glacier surface velocity under steady-state preindustrial conditions. The validation domain predicts grounding-line change rates from the first decade of this century. The application domain predicts glacier mass change in 2100. Figure 3 illustrates these distinct domains. When transitioning between domains, data shifts across domains must be considered. For example, a model trained only on calibration data from recent ice-sheet forcings may fail to predict ice-sheet properties under different future conditions.
@@ 3.3 Quantities of interest
Quantities of interest (QoI) are the model outputs required by users. Their form and scale depend on modeling purpose and application domain. We now discuss key considerations for QoI selection.
Risk assessment requires reproducing only decision-critical QoI. For ice-sheets, these include sea-level rise from mass loss and infrastructure damage costs. Design applications similarly need few QoI to evaluate objectives and constraints, like thermal and structural stresses in aerospace vehicles. Design models need accurate QoI predictions only along optimizer trajectories222For each design iteration the model may still need to be accurate across all uncertain model inputs, while risk assessment models must predict across many conditions. Explanatory modeling demands more extensive QoI sets, such as complete ice-sheet depth and velocity fields for studying calving. Therefore, simple surrogates often suffice for risk assessment and design, but explanatory modeling may require operators or reduced order models.
@@ 3.4 Model conceptualization
Model conceptualization, which follows problem definition, involves selecting model structure based on prior knowledge. While essential to CSE model development [45], this step requires clear identification of the application domain and relevant QoI.
Model structure selection draws on key prior knowledge: conservation laws, system invariances like rotational and translational symmetries. These guide method selection—for example, symplectic time integrators [86] preserve system dynamics properties. Moreover, this knowledge informs and justifies the selection of candidate model structures. A CSE modeler chooses between model types like lumped versus distributed PDE models, and linear versus nonlinear PDEs. The optimal choice depends on application domain, QoI, and available resources. For example, linear PDEs may introduce more error but their lower computational cost enables better error and uncertainty characterization for tasks like optimal design. Similar considerations guide SciML model selection. For example, Gaussian processes excel at predicting scalar QoI with few inputs and limited data, but become intractable for larger datasets without variational inference approximations [61]. In contrast, deep neural networks handle high-dimensional data but require large datasets. The intended use also shapes model structure and training, e.g., optimization applications require controlling derivative errors [10, 76] to ensure convergence [14, 64]. These approximation errors must be understood and quantified where possible.
CSE has a strong history of using prior knowledge to formulate governing equations for complex phenomena and deriving numerical methods that respect important physical properties. However, all models are approximate and the best model must balance usability, reliability, and feasibility [34]. While SciML methods can be usable and feasible, more attention is needed to establish their trustworthiness. In the following two sections, we discuss how CSE V&V can improve the trustworthiness of SciML research.
@@ 4 Verification
Verification increases the trustworthiness of numerical models by demonstrating that the numerical method can adequately solve the equations of the desired mathematical model and the code correctly implements the algorithm. Verification consists of code verification and solution verification, which enhance credibility and trust in the model’s predictions. Code and solution verification are well-established in CSE to reduce algorithmic errors. However, verification for SciML models has received less attention due to the field’s young age and unique challenges. Moreoever, because SciML models heavily rely on data, unlike CSE models, existing CSE verification notions need to be adapted for SciML.
@@ 4.1 Code verification
Code verification ensures that a computer code correctly implements the intended mathematical model. For CSE models, this involves confirming that numerical methods and algorithms are free from programming errors (“bugs”). PDE-based CSE models commonly use the method of manufactured solutions (MMS) to verify code on arbitrarily complex solutions. MMS substitutes a user-provided solution into the governing equations, then differentiates it to obtain the exact forcing function and boundary conditions. These solutions check if the code produces the known theoretical convergence rate as the numerical discretization is refined. If the observed order of convergence is less than theoretical, causes such as software bugs, insufficient mesh refinement, or singularities and discontinuities affecting convergence must be identified.
@@
Code verification for SciML models is important but challenging due to the large role of data and nonconvex numerical optimization. Three main challenges limit code verification for many SciML models. First, while theoretical analysis of SciML models is increasing [93, 94, 78, 57, 55, 53, 54], many models like neural networks do not generally admit known convergence rates outside specific map classes [93, 94, 78, 38], despite their universal approximation properties [40, 20, 57]. Second, generalizable procedures to refine models, such as consistently refining neural-network width and depth as data increases, do not exist. Finally, regardless of data amount and model unknowns, modeling error often plateaus at a much higher level than machine precision due to nonconvex optimization issues like local minima and saddle points [21, 9].
Developing theoretical and algorithmic advances to address the three main challenges limiting code verification can substantially improve the trustworthiness of SciML models. Convergence-based code verification is currently possible only for certain SciML models with theory that bounds approximation errors in terms of model complexity and training data amount, such as operator methods [105], polynomial chaos expansions [16, 112], and Gaussian processes [13].
@@
For SciML models without supporting theory, convergence tests should still be conducted and reported. Studies providing evidence of model convergence engender greater trustworthiness than those that do not, even when the empirically estimated convergence rate cannot be compared to theoretical rates. For example, observing Monte Carlo-type sampling rates in a regime of interest for a fixed overparametrized model can provide intuition into whether the model should be enhanced.
@@
To account for the heavy reliance of SciML models on training data optimization, code verification should be adapted in two ways. First, report errors in the ML model for a given complexity and data amount for different realizations of the training data to quantify the impact of sampling error, which is not present in CSE models. Second, because most SciML algorithms introduce optimization error, conduct verification studies that artificially generate data from a random realization of an ML model, then compare the recovered parameter values with the true parameter values or compare the predictions of the learned and true approximations, or at the very least compare the predictions of the two models. Additionally, quantify the sensitivity of the SciML model error to randomness in the optimization by varying the random seed and initial guess passed to the optimizer (see Section 6.3). All verification tests must employ test data or verification data, independent of the training data, to measure the accuracy of the SciML model.

### v2

@@ 1 Introduction
Scientific machine learning (SciML) integrates machine learning (ML) into scientific workflows to enhance system simulation and analysis, with an emphasis on computational modeling of physical systems. This field emerged from Department of Energy workshops and initiatives starting in 2018, which also identified the need to increase “the scale, rigor, robustness, and reliability of SciML necessary for routine use in science and engineering applications” [5]. The field’s subsequent growth through funding initiatives, conference themes, and high-profile publications stems from its ability to unite ML’s predictive power with the domain knowledge and mathematical rigor of computational science and engineering (CSE). However, this surge in SciML development has outpaced good practices and reporting standards for building trust [55, 41, 95, 104, 99].
@@
Rigorous modeling practices require sustained effort from the scientific community. CSE has established standards over decades for model reporting and assessment, including verification and validation, across biology [66], earth sciences [65, 34], and engineering [82, 71, 76, 1]. These standards help users and decision-makers document limitations, uncertainties, and modeling choices to prevent resource misallocation and maintain model credibility. Similar SciML efforts are “indispensable” [68] but have only recently emerged in scientific disciplines, including materials science [95], environmental science [104], nuclear engineering [96], and aviation [38], alongside more general initiatives [41]. Developers and users of SciML must establish best practices for the development, reporting, and evaluation of new SciML paradigms.
@@ 2.1 Scope
This paper focuses on SciML that enhances computational modeling and simulation of physical systems. Specifically, we address SciML approaches that support computational workflows where a model must be evaluated repeatedly within larger algorithmic procedures. In scientific computing, these are known as “outer-loop” processes because they form an additional computational layer that wraps around the core model solution. Such processes include: (1) explanatory modeling, which searches for insights into complex system behaviors; (2) uncertainty quantification, which infers model parameters from observations and propagates them to predictive distributions [21]; (3) sensitivity analysis, which identifies factors with greatest influence on predictions [75]; (4) experimental design, which predicts which data collections maximize understanding or decision value [74]; and (5) optimal design and control, which optimizes objective functions by adjusting decision variables under constraints [54, Ch. 1]. To support these outer-loop processes, predictive SciML encompasses four main classes of methods, each presenting different computational challenges: surrogate modeling, model discovery, hybrid CSE-SciML modeling, and outer-loop learning.
Surrogate modeling builds low-cost approximations of computationally expensive models for outer-loop analyses. Surrogate models can map model inputs to specific quantities of interest [13, 53]. Surrogate models can also approximate complete solutions to governing equations via data-driven reduced-order modeling [67, 79], operator learning with Gaussian processes [57] and neural networks [51, 46], and physics-informed neural networks (PINNs) [101, 69], which encode physical constraints directly into the neural network training. Finally, surrogate models can provide solutions to differential equations without simulation data [86, 19].
@@
Outer-loop learning leverages ML to predict solutions under varying conditions. Applications include surrogate models for expensive forward simulations in PDE-constrained optimization [97] and Bayesian inverse problems [10]. Reinforcement learning solves PDE-constrained optimization across different conditions [88]. ML methods can also learn solutions to these problems directly. For example, ML models learn Bayesian posterior distributions as marginals of joint distributions for fast online inference with arbitrary observational data [6, 98].
@@ 2.2 Comparing CSE and SciML model development and deployment
Making scientific claims using SciML models involves two steps: model development and model deployment. While SciML models can complement CSE models, their development processes differ, though deployment processes remain similar, as shown in Figure 1. In this paper we discuss the process of quantifying uncertainty in a hybrid SciML-CSE ice-sheet model of the Humboldt Glacier in Greenland, with unknown glacier-landmass friction, to provide an example of how these processes differ. Our example is loosely based on the study by He et al. [28], which exemplifies thorough reporting of a SciML model’s strengths and weaknesses.
@@
CSE model development begins with physics-based governing equations. In our example, nonlinear Stokes equations accurately model shear-thinning behavior for ice sheets, but simpler models like shallow-shelf approximation (SSA) are often used for computational efficiency [28]. The next step implements numerical methods to solve the equations while respecting properties like mass and energy conservation, using techniques such as finite elements, Newton iteration, and time-stepping to compute the coefficients of a fixed ansatz (basis) used to represent the solution. For instance, He and coworkers [28] used continuous piecewise linear elements for thickness and velocity fields with Newton iteration and semi-implicit time-stepping.
@@
In contrast, SciML typically begins by collecting input-output training data from either physical experiments or simulation. For example, He et al. [28] built a DeepONet [51] to map ice-sheet thickness and friction to glacier velocity, using data from an SSA-based CSE model. The second step sets model hyperparameters (e.g., neural-network width and depth). The subsequent steps involve optimizing the model parameters (e.g., neural-network weights and biases) to minimize the difference between model outputs and training data. He et al. [28] used Adam optimization to minimize the mean-squared error between predicted and training velocities. Note that often, hyperparameters must be tuned through formal optimization or informal assessment of model accuracy across different settings. Additionally, some SciML models [86, 19] combine CSE and SciML development steps shown in Figure 1, training without data by minimizing objectives that enforce known governing equations.
@@
The development and deployment of SciML models for specific scientific applications follows distinct phases. When using these models to make scientific claims—–like predicting the Humboldt glacier’s impact on sea level in the future—–deployment involves three key steps: 1) collecting calibration data (distinct from training data) representing the system’s current state; 2) calibrating the model through optimization to match this system-specific data, which includes solving an inverse problem for friction parameters based on recent velocity measurements [35], notably optimizing model inputs rather than SciML parameters like neural-network weights; and 3) running the model under various future conditions or plausible friction fields, such as those obtained from Bayesian calibration.
@@ 2.3 Leveraging an existing CSE framework for trustworthy SciML
As highlighted by Broderick et al. [8], trust in SciML and CSE models can be gained or lost at any step of development and deployment (Figure 1). Building trust requires: 1) demonstrating that the mathematical model adequately meets accuracy goals despite being an approximation; 2) verifying it by comparing theoretical and empirical results to ensure it solves its mathematical formulation correctly; 3) rigorously testing code to reduce software bugs; and 4) documenting data provenance to ensure relevance for future conclusions and decisions.
@@
While V&V practices are well-established in scientific computing, the emergence of SciML brings new concerns that existing guidelines don’t fully address. SciML combines aspects of both scientific computing and ML, presenting unique verification and validation challenges beyond either field. We argue that adapting proven V&V principles can help provide a foundation for building trustworthy SciML models. To this end, we make recommendations in four areas: problem definition, verification, validation, and continuous credibility building (see Figure 2). Problem definition and verification (Sections 3 and 4) expand on the model development steps in Figure 1, while validation and reporting (Sections 5 and 6) detail the deployment steps. Validation is generally most pertinent only to papers that utilize SciML to support a scientific claim for a specific application, but all other components illustrated in Figure 1 are applicable to all SciML papers, including those that present algorithmic advances relevant to a broad spectrum of scientific applications.
Like Broderick et al. [8], we view trustworthiness as a spectrum rather than binary. Following Oberkampf’s predictive capability maturity model [62], the four components in Figure 2 build confidence through continuous testing and refinement as new data emerge. Each study should mature progressively in its development process through these four areas. The following sections identify opportunities for improving SciML practice, along with corresponding recommendations. While, we do not claim that these four components or our recommendations are comprehensive or universally applicable, we aim to stimulate community efforts to develop SciML guidelines and standards for SciML development and deployment.
@@ 3.1 Model purpose
A SciML model’s purpose, as discussed in Section 2.1, dictates all subsequent modeling choices, including required outer-loop processes and essential properties. For example, an explanatory model must simulate all system processes, such as ice-sheet thickness and velocity evolution, while a risk assessment model focuses only on decision-relevant quantities like mass loss under varying emissions scenarios. In contrast, design and control models have different requirements, and the model’s purpose influences the data types and formulations needed for training. Therefore, the model formulation should be chosen based on these problem-specific considerations.
@@ 3.2 Verification, calibration, validation and application domain
Trustworthy model development and deployment require assessing performance across various operational modes or domains [11]. We advocate for evaluating models in four key domains: verification, calibration, validation, and application (see Figure 3), defined by the conditions under which the model operates, including boundary conditions, forcing functions, geometry, and timescales. For ice sheets, examples include surface mass balance, land mass topography, and ocean temperatures.
Each domain typically requires predicting different quantities of interest under varying conditions, with process complexity increasing from verification to application, while the amount of supporting data decreases. For instance, the verification domain predicts the entire state of the ice sheet using simple solutions, while the calibration domain focuses on Humboldt Glacier surface velocity under steady-state conditions, the validation domain predicts grounding-line change rates from the last decade, and the application domain forecasts glacier mass change in 2100. Data shifts between domains must be considered; for example, a model trained solely on recent calibration data may struggle to predict ice-sheet properties under different future conditions.
@@ 3.3 Quantities of interest
Quantities of interest (QoI) are the model outputs required by users, with their form and scale depending on the modeling purpose and application domain. Risk assessment focuses on decision-critical QoIs, such as sea-level rise and infrastructure damage costs for ice sheets, while design applications require fewer QoIs to evaluate objectives and constraints, like thermal and structural stresses in aerospace vehicles. In contrast, Design models need accurate QoI predictions only along optimizer trajectories while risk assessment models must predict across many conditions 222For any iteration of a risk averse design optimization the model may still need to be accurate across all uncertain model inputs. Whereas, explanatory modeling requires a broader range of QoIs, such as complete ice-sheet depth and velocity fields for studying calving, indicating that simple surrogates may suffice for risk assessment and design, but explanatory modeling may necessitate operators or reduced-order models.
@@ 3.4 Model conceptualization
Model conceptualization, following problem definition, involves selecting model structure based on prior knowledge and requires clear identification of the application domain and relevant QoI. Key prior knowledge, such as conservation laws and system invariances, should guide method selection – for example, symplectic time integrators can be used to preserve system dynamics [73].
A CSE modeler chooses between model types, such as lumped versus distributed PDE models, and linear versus nonlinear PDEs, with the optimal depending on the application domain, QoI, and available resources. For example, linear PDEs may introduce more error but their lower computational cost enables better error and uncertainty characterization for tasks like optimal design. Similar considerations guide SciML model selection. For example, Gaussian processes excel at predicting scalar QoI with few inputs and limited data, but become intractable for larger datasets without variational inference approximations [49]. In contrast, deep neural networks handle high-dimensional data but require large datasets. The intended use also shapes model structure and training, e.g., optimization applications require controlling derivative errors [7] to ensure convergence [52].
CSE has a strong history of using prior knowledge to formulate governing equations for complex phenomena and deriving numerical methods that respect important physical properties. However, all models are approximate and the best model must balance usability, reliability, and feasibility [25]. While SciML methods can be usable and feasible, more attention is needed to establish their trustworthiness, which we will address in the following two sections on V&V.
@@ 4 Verification
Verification increases the trustworthiness of numerical models by demonstrating that the numerical method can adequately solve the equations of the desired mathematical model and the code correctly implements the algorithm. Verification consists of code verification and solution verification, which enhance credibility and trust in the model’s predictions. Code and solution verification are well-established in CSE to reduce algorithmic errors. However, verification for SciML models has received less attention due to the field’s young age and unique challenges. Moreover, because SciML models heavily rely on data, unlike CSE models, existing CSE verification notions need to be adapted for SciML.
@@ 4.1 Code verification
Code verification ensures that a computer code correctly implements the intended mathematical model. For CSE models, this involves confirming that numerical methods and algorithms are free from programming errors (“bugs”). PDE-based CSE models commonly use the method of manufactured solutions (MMS) to verify code against known theoretical convergence rates by substituting a user-provided solution into the governing equations. If the observed order of convergence is less than theoretical, potential causes such as software bugs or insufficient mesh refinement must be identified.
@@
Code verification for SciML models is also important but challenging due to the large role of data and nonconvex numerical optimization. Three main challenges limit code verification for many SciML models: while theoretical analysis of SciML models is increasing [81, 64, 44, 43], many models like neural networks do not generally admit known convergence rates outside specific map classes [81, 64, 29]; there are not general procedures for refining models, e.g. changing neural-network width and depth as data increases, and generalization error often plateaus due to nonconvex optimization issues like local minima and saddle points [17].
@@
Addressing these challenges through theoretical and algorithmic advances can enhance the trustworthiness of SciML models. Currently, convergence-based verification is feasible for certain models with established theory, such as operator methods [91], polynomial chaos expansions [13], and Gaussian processes. For models without supporting theory, convergence tests should still be conducted and reported, as evidence of model convergence increases trustworthiness. For example, observing Monte Carlo-type sampling rates in a regime of interest for a fixed overparametrized model can provide intuition into whether the model should be enhanced.
@@
To adapt code verification for SciML models, first, report errors for varying training data realizations to quantify sampling error. Second, conduct verification studies using artificially generated data from a random realization of a SciML model, comparing recovered parameter values or predictions with true values. Additionally, quantify the sensitivity of model error to randomness in optimization by varying the random seed and initial guess (see Section 6.3). All verification tests must use independent test or verification data to measure the accuracy of the SciML model.